OpenStudy (anonymous):

to solve for a variable involving an nth root, raise both sides of the equation to the nth power (y√n)=y f(x)=x√+2

OpenStudy (anonymous):

\[(\sqrt[n]{y})^n =y \] \[\sqrt{x} +2\]